# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

### v3

@@ Appendix D Comprehensive Evaluation across STEM Weight Configurations
Appendix E Syntax Token Coverage in SWEET Watermarking
Table 10 presents an analysis of token selection by SWEET using two measures: (1) the proportion of selected tokens exceeding the entropy threshold among all generated tokens, and (2) the proportion of syntax-related tokens among those selected. We use the optimal settings ($\gamma$ = 0.25, $\delta$ = 3.0, entropy threshold = 0.9)) and evaluate threshold values in the range [0.7, 0.8, 0.9, 1.0, 1.1]. The results show an inverse relationship between the threshold and token coverage: lower thresholds select more tokens overall but include more syntax elements, whereas higher thresholds reduce syntax token inclusion. This pattern shows that SWEET effectively targets high-entropy regions but does not completely avoid syntax-related tokens.
@@ Appendix E Syntax Token Coverage in SWEET Watermarking
Table 12 lists representative examples. Delimiters such as colons (‘:’) and commas (‘,’), essential for code structure, cause parsing errors when modified (e.g., omitting a colon in ‘for i in range(10):’ leads to a failure). Although keywords and operators appear less frequently, their modification can still alter program behavior, for example, replacing ‘True’ with ‘False’ or ‘+’ with ‘-’. Minimizing the inclusion of such syntax-critical tokens can make syntax-aware watermarking more reliable and less disruptive to functionality.
@@ Appendix F Comparison with Non-watermarked Code
We evaluate the functional impact of watermarking by comparing the correctness of code generated with and without watermark insertion. Non-watermarked code defines the upper bound of functional performance. STONE achieves scores nearly identical to this bound, showing that watermark insertion has minimal influence on code execution. Among all watermarking methods, STONE reduces correctness the least, which indicates that it embeds provenance information effectively while preserving the original program behavior.

### v4

@@ Appendix D Comprehensive Evaluation across STEM Weight Configurations
Appendix E Syntax Token Coverage in SWEET
Table 11 presents an analysis of token selection by SWEET using two measures: (1) the proportion of selected tokens exceeding the entropy threshold among all generated tokens, and (2) the proportion of syntax-related tokens among those selected. We use the optimal settings ($\gamma=0.25$, $\delta=3.0$, entropy threshold =0.9=0.9) and evaluate threshold values in the range $\{0.7,0.8,0.9,1.0,1.1\}$. The results show an inverse relationship between the threshold and token coverage: lower thresholds select more tokens overall but include more syntax elements, whereas higher thresholds reduce syntax token inclusion. This pattern shows that SWEET effectively targets high-entropy regions but does not completely avoid syntax-related tokens.
@@ Appendix E Syntax Token Coverage in SWEET
Table 12 lists representative examples. Delimiters such as colons (‘:’) and commas (‘,’), essential for code structure, can cause parsing errors when modified (e.g., omitting a colon in ‘for i in range(10):’ leads to a failure). Although keywords and operators appear less frequently, their modification can still alter program behavior, for example, replacing ‘True’ with ‘False’ or ‘+’ with ‘-’. Minimizing the inclusion of such syntax-critical tokens can make syntax-aware watermarking more reliable and less disruptive to functionality.
@@ Appendix F Comparison with Non-watermarked Code
We evaluate the functional impact of watermarking by comparing the correctness of code generated with and without watermark insertion. Non-watermarked code provides an empirical upper bound in our evaluation setting. STONE achieves scores nearly identical to this bound, showing that watermark insertion has minimal influence on code execution. Among all watermarking methods, STONE reduces correctness the least, which indicates that it embeds provenance information effectively while preserving the original program behavior.